zarkam21:

Help please on these 2 questions

zarkam21:

1 attachment
Vocaloid:

 you'll have to complete the square on the x-terms and y-terms, are you familiar with how to do this?

zarkam21:

NO :/

Vocaloid:

start by isolating the x terms x^2 - 8x now, take the b-coefficient (-8), divide by 2, then square it - what do you get?

zarkam21:

-8/2=-4 2i

Vocaloid:

square it not take the square root

zarkam21:

-4^2

Vocaloid:

be careful with parentheses (-4)^2 = ?

zarkam21:

16

Vocaloid:

good, so we add 16 to both sides of the equation now that we have (x^2 - 8x + 16) factor this into a squared binomial

zarkam21:

(x-4)^2

Vocaloid:

good, now we repeat this w/ the y terms y^2 +10y again, take the b-term, divide it by 2, square it, add it to both sides of the eqn, then factor

zarkam21:

10/2=5 5^2

zarkam21:

25

Vocaloid:

good, then factor y^2 + 10y + 25

zarkam21:

(y+5)^2

Vocaloid:

good, so far we have (x-4)^2 + (y+5)^2 = 16 + 25 = 41 now, divide both sides by 41 to get it into standard form (x-4)^2 / 41 + (y+5)^2 / 41 = 1 what conic is this? take a look at the thing I posted at the beginning

zarkam21:

ellipse

Vocaloid:

awesome now, looking at the ellipse equation where the center is (h,k) what's the coordinates of the center?

Vocaloid:

(not a hyperbola, we have a + sign not a - sign in the middle)

zarkam21:

(4,-5)

Vocaloid:

other way around with the signs

zarkam21:

okay so (-4,5)

Vocaloid:

anyway for #2 try completing the square w/ the y terms and re-arranging the equation for x

Vocaloid:

y^2 +4y, complete the square

zarkam21:

1 attachment
Vocaloid:

the second equation is not an ellipse or hyperbola (no x^2 term) so it has to be written in the form x = a(y-k)^2 + h

zarkam21:

the -8(x-2) one?